Prev Next

SW&IT Meeting Minutes 2021-10-11

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20211004
LR20211004 - Bec, Matthieu 40.0
LR20211004 - Mechtley, Brandon 40.0
LR20211004 - Eychaner, Glenn 40.0
LR20211004 - Edmeades, David 40.0
LR20211004 - Grenz, Paul 20.0
LR20211004 - Hooper, Stephen 80.0
LR20211004 - Zhang, Yang 40.0
LR20211004 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • house for sale spam on sciops

Team reports

MB

  • Last week
    • OC, Management
    • ATT PM#8, SX AO restart meeting
    • learning some more about AO s/w, join discussions and scheming on #ao-sw
    • more AO/dms discussions with Al
  • This week
    • git/github "policy" (= philosophy slides) for new staff
    • AO s/w pause-resume-ramp assistance as needed
    • review TCS punch-list from JHill with PGrenz
    • look into broken imonnit (stopped Sep 27)

YZ

  • Previous:
    • Continue some work on MODS WCS program (revisions, github commit, brief discussion with Olga).
    • TMS meeting, discussios, and software work.
    • Worked with Paul on Argos swingarm tests and debugging.
  • Next:
    • Maybe check with archive team on image notification questions?
    • Work with Olga on verifying the MODS WCS program.
    • Other misc LBC/MODS stuffs.
    • TMS meeting and other work to continue.
    • Continue to work with Paul on Argos swingarms tests and debugging.

PSG

  • Last week:
    • Testing Swing Arms again.
    • Strange behavior of D425 not acting on commands, eventually it worked though.
    • HL2000 would not allow connection, we may need to re-examine the design.
  • This week:
    • Research why D425 did not respond at first.
    • Determine the best path forward with the HL2000
    • Discuss additional work suggested by John Hill.

DE

  • Backup server work
    • Got it into a bad state re: updates, learned slipstream, will reinstall with fresh W10 image that includes many updates + 10GbE NIC driver
  • Windows server license, need to talk to Ruben, who is back from vacation today
  • VPN account; reiterate impressing on people that we need more notice from the LBT sponsors when they request accounts
  • House spam wasn't on sciops, it was from sciops. I don't think marking this as spam is going to cause false positives, especially with the keywords involved.
    • If this becomes a problem we can tighten the group restrictions, but they are open for a reason. We'll just have to gauge the acceptable level and hope that the cyclical nature of spammers moves to a different tactic.
    • Re-added Laura to sciops
  • Laser cutter computer situation cleared up with Robert.
  • Mountain Synology volume filling
    • Cristina said she'd look into the archive self-cleaning, but that hasn't happened yet.
    • Down to ~1TB, maybe 2 days
    • Should we start manually clearing? Ping Cristina one more time?
  • Met with JP re: Google Sites transition
    • I opened a ticket with Google to help her get answers to some specific questions
    • She will start tests and see what we're in for
  • Dell techdirect followup; I think I'm confusing their system with LBT vs. UA stuff. I'll talk to a human.

GE

  • Last week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Secured a copy of the original Aladin v10 source code from CDS.
      • Most changes for Aladin for LUCI were incorporated into Aladin, except some last minute synch changes made in response to IT5815.
      • These changes may be responsible for some of the behavior issues such as IT7952, because they allow scripts to execute asynchronously.
  • This week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Read GEIRS manual and look into using GEIRS to save a summed image along with each cube. This would prevent large cubes from freezing Aladin.

BM

  • Last week:
    • AO modified pause/resume routines
    • Started looking at how to continue housekeeper/fastdiagnostic logging when switchbcu communication times out.
  • This week:
    • Tuesday (through Thursday) test new pause/resume routines--need to prepare on soul-sxwfs
    • Other AO issues: logging during switchbcu timeout, logging field width fix
    • Overcurrent investigation
    • AO data migration

SH

  • Mountain calls
  • JW Laptop
  • SOUL Linux
  • UA Risk Management

next:

  • SOUL Linux
  • Workshop
  • In to work today
Topic revision: r10 - 11 Oct 2021, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback